OpenStudy (anonymous):

Is anyone good at slope intercept form? I have two equations that I need help converting and I already know a little bit I just have some trouble. Thanks

OpenStudy (anonymous):

slope intercept form- y=mx+b

OpenStudy (anonymous):

equation 1: 5x+y=9

OpenStudy (anonymous):

Equation 2: 3x+2y=4

OpenStudy (fibonaccichick666):

ok so, how do you bring you 5x to the other side?

OpenStudy (fibonaccichick666):

how do you undo the addition of 5x?

OpenStudy (anonymous):

do I subtract 5 from 5 and 9?

OpenStudy (fibonaccichick666):

not quite, you have \(5\color\red{x}\) not 5,

OpenStudy (anonymous):

so how do I do it then?

OpenStudy (anonymous):

I'm not sure what to subtract first

OpenStudy (fibonaccichick666):

por que no los dos?

OpenStudy (fibonaccichick666):

if you just had x+y=2, how would you move the x to the other side?

OpenStudy (anonymous):

oh ok so if I subtract 5x from both sides I am left with y=9-5x?

OpenStudy (fibonaccichick666):

yep

OpenStudy (anonymous):

ok then do I divide both sides by 9?

OpenStudy (anonymous):

??

OpenStudy (fibonaccichick666):

no why would you? you are done

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

and for equation B?

OpenStudy (fibonaccichick666):

you can rearrange if you want

OpenStudy (fibonaccichick666):

B is the same idea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (fibonaccichick666):

actually, rearrange so you have exactly y=mx+b style currently you have y=b-mx

OpenStudy (fibonaccichick666):

so what is another way to write subtraction?